# Computing Generalized Ranks of Persistence Modules via Unfolding to Zigzag Modules

### v3

@@ 7 Special case of degree-$d$ homology for $d$-complexes
First assume that $p_{i}$ and $p_{j}$ are immediate points in $P_{ZZ}$ and without loss of generality let $p_{i}\rightarrow p_{j}$. According to the definition of representatives, we must have the homology classes $[z_{p_{i}}]$ and $[z_{p_{j}}]$ homologous in $K_{p_{j}}$. Since $K_{p_{j}}$ is a $d$-complex, $[z_{p_{i}}]=[z_{p_{j}}]$ only if $z_{p_{i}}=z_{p_{j}}$ as chains. It follows by transitivity that this is true even if $p_{i}$ and $p_{j}$ are not immediate. ∎
It follows from the above proposition that a full interval module in a direct decomposition of $\mathbb{M}_{ZZ}$ is foldable. Next proposition says that even the complement $\overline{\mathbb{I}}$ is foldable. Then, applying Theorem 3.1(1), we can claim that ${\sf{rk}}(\mathbb{M})$ is equal to the number of full interval modules in $\mathbb{M}_{ZZ}$.
@@
Let $\mathbb{I}$ be any full interval module in a direct decomposition of $\mathbb{M}_{ZZ}$ where $\mathbb{M}_{ZZ}$ is constructed as in Proposition 7.1. Then, both $\mathbb{I}$ and $\overline{\mathbb{I}}$ are foldable.
@@
$\mathbb{I}$ is foldable due to Proposition 7.1. Let $p,p^{\prime}\in P_{ZZ}$ be any two points with $s(p)=s(p^{\prime})$. We need to show that $\overline{\mathbb{I}}(p)=\overline{\mathbb{I}}(p^{\prime})$.
For the full module $\mathbb{I}$, there is a single $d$-cycle that forms a fixed representative at each point in $P_{ZZ}$ (Proposition 7.1). Let $z$ be such a $d$-cycle and $\sigma$ be any $d$-simplex in $z$. Delete $\sigma$ from the complex $K_{p}$ for every $p\in P_{ZZ}$. Let $\mathbb{M}_{ZZ}^{-}$ denote the $P_{ZZ}$-module induced by the homology functor in degree $d$ on the zigzag filtration $F_{ZZ}^{-}$ obtained from the original filtration $F_{ZZ}$ by deleting the simplex $\sigma$ everywhere. It is easy to verify that $\mathbb{M}_{ZZ}^{-}$ is $s$-foldable because $K_{p}\setminus\sigma=K_{p^{\prime}}\setminus\sigma$ for every $p,p^{\prime}\in P_{ZZ}$ with $s(p)=s(p^{\prime})$. A direct decomposition of $\mathbb{M}_{ZZ}$ is obtained from a direct decomposition of $\mathbb{M}_{ZZ}^{-}$ by adding a full bar with the representative $z$. Therefore, $\overline{\mathbb{I}}$ is equal to $\mathbb{M}_{ZZ}^{-}$ and thus foldable. ∎
@@
By Proposition 7.2, $\kappa(\mathcal{D})=\tau(\mathcal{D})$ for any direct decomposition $\mathcal{D}$ of $\mathbb{M}_{ZZ}$. Then, it follows from Proposition 6.2 that ${\sf{rk}}(\mathbb{M})=\kappa(\mathcal{D})$, which is exactly equal to the number of full interval modules in any direct decomposition of $\mathbb{M}_{ZZ}$. Therefore, ${\sf{rk}}(\mathbb{M})$ can be simply obtained by computing the zigzag barcode of $\mathbb{M}_{ZZ}$ (no need of computing the representatives). This can be done in $O(t^{\omega})$ time with the fast zigzag algorithm [14]. ∎

### v4

@@ 7 Special case of degree-$d$ homology for $d$-complexes
First assume that $p_{i}$ and $p_{j}$ are immediate points in $P_{ZZ}$ and without loss of generality let $p_{i}\rightarrow p_{j}$. According to the definition of representatives, we must have the homology classes $z_{p_{i}}]$ and $z_{p_{j}}]$ homologous in $K_{p_{j}}$. Since $K_{p_{j}}$ is a $d$-complex, $z_{p_{i}}]=[z_{p_{j}}]$ only if $z_{p_{i}}=z_{p_{j}}$ as chains. It follows by transitivity that this is true even if $p_{i}$ and $p_{j}$ are not immediate. ∎
It follows from the above proposition that a full interval module $\mathbb{I}$ in a direct decomposition of $\mathbb{M}_{ZZ}$ is foldable. Next proposition says that even the complement $\overline{\mathbb{I}}$ is invertible. Then, applying Theorem 3.1(3), we can claim that $\mathrm{Fld}_{s}(\mathbb{I})$ is a full interval summand of $\mathbb{M}$.
@@
Let $\mathbb{I}$ be any full interval module in a direct decomposition of $\mathbb{M}_{ZZ}$ where $\mathbb{M}_{ZZ}$ is constructed as in Proposition 7.1. Then, $\mathbb{I}$ is foldable and $\overline{\mathbb{I}}$ is invertible.
@@
$\mathbb{I}$ is foldable due to Proposition 7.1. We need to show that $\overline{\mathbb{I}}$ is invertible.
For the full module $\mathbb{I}$, there is a single $d$-cycle that forms a fixed representative at each point in $P_{ZZ}$ (Proposition 7.1). Let $z$ be such a $d$-cycle and $\sigma$ be any $d$-simplex in $z$. Delete $\sigma$ from the complex $K_{p}$ for every $p\in P_{ZZ}$. Let $\mathbb{M}_{ZZ}^{-}$ denote the $P_{ZZ}$-module induced by the homology functor in degree $d$ on the zigzag filtration $F_{ZZ}^{-}$ obtained from the original filtration $F_{ZZ}$ by deleting the simplex $\sigma$ everywhere. It is easy to verify that $\mathbb{M}_{ZZ}^{-}$ is $s$-foldable because $K_{p}\setminus\sigma=K_{p^{\prime}}\setminus\sigma$ for every $p,p^{\prime}\in P_{ZZ}$ with $s(p)=s(p^{\prime})$. A direct decomposition of $\mathbb{M}_{ZZ}$ is obtained from a direct decomposition of $\mathbb{M}_{ZZ}^{-}$ by adding a full bar with the representative $z$ or equivalently the summand $\mathbb{I}$. It follows that $\overline{\mathbb{I}}$ is invertible. ∎
@@
By Proposition 7.2 and Theorem 3.1(3), we get that $\kappa(\mathcal{D})=\tau(\mathcal{D})$ for any direct decomposition $\mathcal{D}$ of $\mathbb{M}_{ZZ}$. Then, it follows from Proposition 6.2 that ${\sf{rk}}(\mathbb{M})=\kappa(\mathcal{D})$, which is exactly equal to the number of full interval modules in any direct decomposition of $\mathbb{M}_{ZZ}$. Therefore, ${\sf{rk}}(\mathbb{M})$ can be simply obtained by computing the zigzag barcode of $\mathbb{M}_{ZZ}$ (no need of computing the representatives). This can be done in $O(t^{\omega})$ time with the fast zigzag algorithm [14]. ∎
@@ 8 Conclusions and discussions
Acknowledgment.
We thank Iason Papadopoulos of University of Bremen for useful comments on the paper.